1. In a ramp experiment, a trolley of mass 0.5 kg is pushed at a velocity of 0.3m/s. It collides with a second stationary trolley of mass 1.0kg and sticks to it. After collision, both trolleys move off with a velocity of 0.1m/s. Has momentum been conserved?
1
Expert's answer
2021-05-21T10:05:33-0400

According to the law of conservation of momentum, we have:


"p_i=p_f,""m_1v_{1,i}+m_2v_{2,i}=(m_1+m_2)v_f,""0.5\ kg\cdot0.3\ \dfrac{m}{s}+0=(0.5\ kg+1\ kg)\cdot0.1\ \dfrac{m}{s},""0.15\ \dfrac{kg\cdot m}{s}=0.15\ \dfrac{kg\cdot m}{s}."

Therefore, momentum has been conserved.